OpenStudy (anonymous):

I forget how to do this... n^12 / n^4 = __________

OpenStudy (anonymous):

power rules. 12-4=8 n^12/n^4 = n^8

OpenStudy (anonymous):

oh ok, thanks haha i remembered you either subtracted or divided :)